The legnths of two adjacent sides of a parallelogram are 6 and 15. If the measure of an included angle is 60 degrees. What is the legnth of the shorter diagnol of the parallelogram?
Answer by ankor@dixie-net.com(22740)

The lengths of two adjacent sides of a parallelogram are 6 and 15. If the measure of an included angle is 60 degrees. What is the length of the shorter diagonal of the parallelogram?
:
We know the short diagonal is opposite the smaller angle, in this case it's 60 deg
:
Draw the triangle, Let A = 60 degrees, the diagonal = a, b = 15, c = 6
:
We have side-angle-side here which lends itself to the law of cosines:
a^2 = b^2 + c^2 - 2(b*c)cos(A)
:
a^2 = 15^2 + 6^2 - 2(15*6)Cos(60)
:
a^2 = 225 + 36 - 2(90)(.5)
:
a^2 = 261 - 90
:
a = SqRt(171)
:
a = 13.1 is the length of the shorter diagonal
;